# On the exact self-similar finite-time blowup of the Hou–Luo model with smooth profiles

### v2

@@ Abstract.
1 dhuang@math.pku.edu.cn
2 jnwdyi@pku.edu.cn
School of Mathematical Sciences, Peking University
@@ 1. Introduction
where the profiles $\widetilde{\Omega}(\cdot,t),\widetilde{\Theta}(\cdot,t)$ and the scaling factors $\tilde{c}_{\omega}(t),\tilde{c}_{\theta}(t),\tilde{c}_{l}(t)$ all depend on time and will converge to some non-trivial steady state as $t\rightarrow T$. In particular, $\tilde{c}_{\omega}(t)$ must converge to $-1$. In their computer-assisted framework, Chen, Hou, and Huang [3] proved the existence of a family of asymptotically self-similar finite-time blowups of the form (1.4) by showing the nonlinear quasi-stability of the corresponding dynamic rescaling equations around an approximate steady state constructed by numerical methods. The term “a family of” means that the initial state of (1.4) can be taken arbitrarily as long as it falls in a small energy-norm ball centered at the approximate steady state (up to rescaling). Furthermore, they also used a limit argument to show that, near the approximate steady state, there lies a true stable steady state that corresponds to exact self-similar profiles, though they could not provide accurate descriptions of these profiles. Nevertheless, their result gives a very accurate estimate of the spacial scaling: $|c_{l}-\bar{c}_{l}|\leq 6\times 10^{-5}$, where $\bar{c}_{l}=2.99870$ is a numerically computed constant corresponding to the approximate steady state, and the bound $6\times 10^{-5}$ results from computer-assisted estimates. It is interesting that $c_{l}$ is extremely close to but strictly smaller than $3$.
@@
The 1D Hou–Luo model (1.1) on the real line admits an exact self-similar finite-time blowup solution of the form (1.3) with $c_{\omega}=-1$, $c_{l}\in(2\,,4.53)$, $c_{\theta}=2c_{\omega}+c_{l}$, and a pair of profiles $\Omega,\Theta$ that satisfy the following:
@@
    The functions $f(x)=\Omega(x)/x$ and $m(x)=\Theta^{\prime}(x)/x$ are both decreasing in $x$ and convex in $x^{2}$ on $[0,+\infty)$; that is, $f^{\prime}(x),m^{\prime}(x)<0$ and $(f^{\prime}(x)/x)^{\prime},(m^{\prime}(x)/x)^{\prime}\geq 0$ for all $x>0$.
@@
Let us remark on our result. The existence of the exact self-similar profiles is proved via a nonlinear fixed-point method. More precisely, we first construct two nonlinear nonlocal maps $\bm{M}$ and $\bm{R}$ over a suitable function $D$ such that, if $f\in\mathbb{D}$ is a fixed point of $\bm{R}$, i.e. $f=\bm{R}(f)$, then $\Omega(x)=xf(x)$ and $\Theta^{\prime}(x)=(c_{l}/2)x\bm{M}(f)(x)$ are a pair of exact self-similar profiles of the HL model (1.1) with $c_{\omega},c_{\theta},c_{l}$ given explicitly in terms of integrals of $f$. We then prove the existence of a fixed point of $\bm{R}$ using the Schauder fixed-point theorem. A key observation in our proof is that the map $\bm{R}$ preserves the properties that $f(x)$ is non-increasing in $x$ and convex in $x^{2}$ for $x\geq 0$, which will be frequently used in our arguments. Furthermore, based on the fixed-point relation $f=\bm{R}(f)$, we are able to determine the regularity of $f,\bm{M}(f)$ and their far-field decay rates, which then transfer to desired properties of the corresponding self-similar profiles $\Omega$ and $\Theta$.
This proof strategy is modified from the fixed-point method developed in a recent work by the same authors [9], where we proved the existence of exact self-similar finite-time blowups of the a-parameterized family of the generalized Constantin–Lax–Majda (gCLM) equation [8, 6, 14],
$\omega_{t}+au\omega_{x}=u_{x}\omega,\quad u_{x}=\bm{H}(\omega),$
for all $a\leq 1$. This equation is also a 1D model for the vorticity formulation of the 3D incompressible Euler equations, with a parameter $a$ that controls the competition between advection and vortex stretching. The main difficulty of modifying our fixed-point method from the gCLM case to the HL case lies in that the gCLM model is an equation of one scalar $\omega$, while the HL model is a coupled system of two scalars $\omega,\theta$. Surprisingly, though the nonlinear map $\bm{R}$ for the HL system is formally much more complicated than the one for the gCLM equation, it still enjoys the critical property that it preserves the previously mentioned monotonicity and convexity of $f\in\mathbb{D}$. From a more essential perspective, this nice property of $\bm{R}$ owes to the kernel structure of the Hilbert transform $\bm{H}$. However, we believe that this fixed-point framework can be further generalized to prove the existence of an exact self-similar finite-time blowup of the 2D Boussinesq equations. This shall be our next step in this line of research.
@@
One unsatisfying thing about our result is the crude estimate of $c_{l}$. Based on our fixed-point method, we can only show that $2<c_{l}\leq 2(k+1)/(k-1)\approx 4.5298$, where $k=1+\sqrt{10}/2$. This is of course much worse than the estimate $2.99870\pm 6\times 10^{-5}$ obtained in [3]. From this comparison, one sees clearly how rigorous computer-assisted estimates can outperform pure analytic estimates in providing sharp bounds.
@@
Finally, we remark that our work does not prove the uniqueness (up to rescaling) of self-similar profiles for the HL model (1.1). Hence, we cannot conclude that the self-similar profiles $\Omega,\Theta$ obtained by our fixed-point method are identical (under proper rescaling) to those obtained via the computer-assisted proof. Nevertheless, it is likely that the exact self-similar solution is unique provided that $c_{\omega}=-1$ and $\Omega^{\prime}(0)=1$. In fact, we can numerically compute the fixed point $f=\bm{R}(f)$ by an iterative scheme and compare it to the one obtained numerically in [3] by solving the dynamic rescaling equation, and we see that they match perfectly well up to only scheme errors.
@@ 2.1. Self-similar profiles
$\begin{split}&-(T-t)^{c_{\omega}-1}c_{\omega}\Omega+(T-t)^{c_{\omega}-1}c_{l}X\Omega_{X}+(T-t)^{2c_{\omega}}U\Omega_{X}=(T-t)^{c_{\theta}-c_{l}}\Theta_{X},\\ &-(T-t)^{c_{\theta}-1}c_{\theta}\Theta+(T-t)^{c_{\theta}-1}c_{l}X\Theta_{X}+(T-t)^{c_{\theta}+c_{\omega}}U\Theta_{X}=0,\\ &U_{X}=\bm{H}(\Omega),\quad U(0)=0,\end{split}$
@@
It is worth mentioning that Chen, Hou, and Huang [3] studied the dynamic rescaling equations of the HL model (see [3, Equation (2.1)]),
@@
Finally, we remark that it is possible to find odd nontrivial solutions of equations (2.1) that is more degenerate at $x=0$ in the sense that $\omega^{\prime}(0)=v^{\prime}(0)=0$ while $\omega^{(n)}(0)\neq 0,v^{(n)}(0)\neq 0$ for some odd integer $n>1$. Numerical evidence of the existence of these degenerate solutions has been reported in the thesis of Liu [12].
@@ 2.2. Reformulation of the problem
From the second equation of (2.8), we easily get
@@
meaning that $m$ can be explicitly determined by $g$ and hence by $f$. Next, we rearrange the first equation of (2.8) to get
@@
[…]
@@
[…]
From the derivation above, it is apparent that a solution of the integral equations (2.9) corresponds to a solution of our main equations (2.1) via the change of variables (2.6), (2.7).
@@ 3. Existence of solution by a fixed-point method
Our goal of this section is to show that the nonlinear integral equations (2.9) admit a non-trivial solution. As we can see, the system (2.9) naturally defines a fixed-point problem of a nonlinear nonlocal map. To prove the existence of a suitable fixed-point solution, we need to construct some appropriate function set in a Banach function space on which we can establish continuity and compactness of this nonlinear map, and then we use the Schauder fixed-point theorem.
@@ 3.1. A fixed-point problem
$\mathbb{D}=\Big{\{}f\in\mathbb{V}:\quad$ $f(x)\geq 0,\ f(0)=1,\ \text{$m_{1}(x)\leq f(x)\leq 1$ for all $x$},$
$f^{\prime}_{-}(1)\leq-\eta,\ \text{$f(x)$ is non-increasing on $[0,+\infty)$ },\ \text{$f(\sqrt{s})$ is convex in $s$},$
@@
The upper bound of $f$ in (3.2) follows from the assumptions that $f(\sqrt{s})$ is convex in $s$, $f^{\prime}_{-}(1)\leq-\eta$, and $f(x)$ is non-increasing on $[0,+\infty)$.
We remark that, though a function $f\in\mathbb{D}$ is not required to be differentiable, the one-sided derivatives $f^{\prime}_{-}(x)$ and $f^{\prime}_{+}(x)$ are both well defined at every point $x$ by the convexity of $f(\sqrt{s})$ in $s$. In what follows, we will abuse notation and simply use $f^{\prime}(x)$ for $f^{\prime}_{-}(x)$ and $f^{\prime}_{+}(x)$ in both weak sense and strong sense. For example, when we write $f^{\prime}(x)\leq C$, we mean $f^{\prime}_{-}(x)\leq C$ and $f^{\prime}_{+}(x)\leq C$ at the same time. In this context, the non-increasing property of $f$ on $[0,+\infty)$ can be represented as $f^{\prime}(x)\leq 0$ for $x\geq 0$.
@@
Comparing this definition with (2.2), one finds that
@@
Finally, letting
@@
We now explain the design of the set $\mathbb{D}$. The guiding idea is to make $\mathbb{D}$ as specific as possible while keeping it compact in the $L^{\infty}_{\rho}$-norm and closed under the map $\bm{R}$, so that we can apply the Schauder fixed-point theorem. The compactness of $\mathbb{D}$ is relatively easy, while its closedness under $\bm{R}$ takes most of the work. Firstly, the monotonicity in $x$ and the convexity in $x^{2}$, which provide powerful controls on the functions in $\mathbb{D}$, are magically preserved by the map $\bm{R}$. This is essentially due to the nice properties of the kernel of the linear map $\bm{T}$. Secondly, the monotonicity, the convexity, and some calculations of $\bm{T}$ will lead to the control $m_{1}(x)\leq\bm{R}(f)\leq 1$. Hence, we can impose this condition on $\mathbb{D}$ as well. Thirdly, we need $\mathbb{D}\subset L^{1}(\mathbb{R})$ in order for $b(f)$ to be well defined on $\mathbb{D}$, which explains the condition $f(x)\lesssim|x|^{-1-\delta_{1}}$. However, to make sure this tail bound is not compromised by $\bm{R}$, we need the weaker estimate $f(x)\lesssim|x|^{-\delta_{0}}$, and to further pass this weaker estimate to $\bm{R}(f)$ we need to use the condition $f^{\prime}_{-}(1)\leq-\eta$. Also, $f^{\prime}_{-}(1)\leq-\eta$ leads to the upper bound in (3.2), which further implies $c(f)>0$, so that $\bm{G}(f)$ and $d(f)$ are both well-defined. Finally, the estimate $\bm{R}(f)^{\prime}(1)\leq-\eta$ can be derived from $f^{\prime}_{-}(1)\leq-\eta$ and the convexity of $f(x)$ in $x^{2}$, therefore closing the loop.

### v3

# Exact self-similar finite-time blowup of the Hou–Luo model with smooth profiles

@@ Abstract.
1 School of Mathematical Sciences, Peking University. Email: dhuang@math.pku.edu.cn
2 Computing and Mathematical Sciences, California Institute of Technology. Email: xqin2@caltech.edu.
3 School of Mathematical Sciences, Peking University. Email: wangxiuyuan@stu.pku.edu.cn
4 School of Mathematical Sciences, Peking University. Email: jnwdyi@pku.edu.cn
@@ 1. Introduction
where the profiles $\widetilde{\Omega}(\cdot,t),\widetilde{\Theta}(\cdot,t)$ and the scaling factors $\tilde{c}_{\omega}(t),\tilde{c}_{\theta}(t),\tilde{c}_{l}(t)$ all depend on time and will converge to some non-trivial steady state as $t\rightarrow T$. In particular, $\tilde{c}_{\omega}(t)$ must converge to 1-1. In their computer-assisted framework, Chen, Hou, and Huang [7] proved the existence of a family of asymptotically self-similar finite-time blowups of the form (1.4) by showing the nonlinear quasi-stability of the corresponding dynamic rescaling equations around an approximate steady state constructed by numerical methods. The term “a family of” means that the initial state of (1.4) can be taken arbitrarily as long as it falls in a small energy-norm ball centered at the approximate steady state (up to rescaling). Furthermore, they also used a limit argument to show that, near the approximate steady state, there lies a true stable steady state that corresponds to exact self-similar profiles, though they could not provide accurate descriptions of these profiles. Nevertheless, their result gives a very accurate estimate of the spacial scaling: $|c_{l}-\bar{c}_{l}|\leq 6\times 10^{-5}$, where $\bar{c}_{l}=2.99870$ is a numerically computed constant corresponding to the approximate steady state, and the bound $6\times 10^{-5}$ results from computer-assisted estimates. It is curious that $c_{l}$ is extremely close to but strictly smaller than $3$.
@@
The 1D Hou–Luo model (1.1) on the real line admits an exact self-similar finite-time blowup solution of the form (1.3) with $c_{\omega}=-1$, $c_{l}\in(2\,,4.53)$, $c_{\theta}=2c_{\omega}+c_{l}$, and a pair of smooth profiles $\Omega,\Theta$ that satisfy the following:
@@
    The functions $f(x):=\Omega(x)/x$ and $m(x):=\Theta^{\prime}(x)/x$ are both non-increasing in $x$ for $x\geq 0$. Moreover, $f(\sqrt{s})$ and $m(\sqrt{s})$ are both convex in $s$ for $s\geq 0$. More precisely, $f^{\prime}(x),m^{\prime}(x)\leq 0$ and $(f^{\prime}(x)/x)^{\prime},(m^{\prime}(x)/x)^{\prime}\geq 0$ for all $x>0$.
@@
Let us remark on our result. The existence of the exact self-similar profiles is proved via a nonlinear fixed-point method. More precisely, we first construct two nonlinear nonlocal maps $\bm{M}$ and $\bm{R}$ over a suitable function $D$ such that, if $f\in\mathbb{D}$ is a fixed point of $\bm{R}$, i.e. $f=\bm{R}(f)$, then $\Omega(x)=xf(x)$ and $\Theta^{\prime}(x)=(c_{l}/2)x\bm{M}(f)(x)$ are a pair of exact self-similar profiles of the HL model (1.1) with $c_{\omega},c_{\theta},c_{l}$ given explicitly in terms of integrals of $f$. We then prove the existence of a fixed point of $\bm{R}$ using the Schauder fixed-point theorem. A key observation in our proof is that the map $\bm{R}$ preserves the properties that $f(x)$ is non-increasing in $x$ for $x\geq 0$ and $f(\sqrt{s})$ is convex in $s$ for $s\geq 0$, which will be frequently used in our arguments. Furthermore, based on the fixed-point relation $f=\bm{R}(f)$, we are able to determine the regularity of $f,\bm{M}(f)$ and their far-field decay rates, which then transfer to desired properties of the corresponding self-similar profiles $\Omega$ and $\Theta$.
@@
One unsatisfying thing about our result is the crude estimate of $c_{l}$. Based on our fixed-point method, we can only show that $2<c_{l}\leq 2(k+1)/(k-1)\approx 4.5298$, where $k=1+\sqrt{10}/2$. This is of course much worse than the rigorous computer-assisted estimate $|c_{l}-2.99870|\leq 6\times 10^{-5}$ obtained in [7]. From this comparison, one sees clearly how rigorous computer-assisted estimates can outperform pure analytic estimates in providing sharp bounds.
@@
Finally, we remark that our work does not prove the uniqueness (up to rescaling) of self-similar profiles for the HL model (1.1). Hence, we cannot conclude that the self-similar profiles $\Omega,\Theta$ obtained by our fixed-point method are identical (under proper rescaling) to those obtained in [7] via the computer-assisted proof. Nevertheless, it is likely that the exact self-similar solution is unique provided that $c_{\omega}=-1$ and $\Omega^{\prime}(0)=1$. In fact, we can numerically compute the fixed point $f=\bm{R}(f)$ by an iterative scheme and compare it to the one obtained numerically in [7] by solving the dynamic rescaling equation, and we see that they match perfectly well up to only scheme errors.
Our proof strategy is modified from the fixed-point method developed in a recent work by the same authors [16], where we proved the existence of exact self-similar finite-time blowups of the a-parameterized family of the generalized Constantin–Lax–Majda (gCLM) equation [23],
$\omega_{t}+au\omega_{x}=u_{x}\omega,\quad u_{x}=\bm{H}(\omega),$ (1.5)
for all $a\leq 1$. This equation is also a 1D model for the vorticity formulation of the 3D incompressible Euler equations, with a parameter $a$ that controls the competition between advection and vortex stretching. The case $a=0$ is the original CLM model [10], the case $a=1$ is also known as the De Gregorio model [12], and the case a=1a=-1 corresponds to the Córdoba–Córdoba–Fontelos model [2]. While the HL model represents a relatively recent development in this field, the gCLM model has been studied extensively for decades. We only mention below some most relevant results on finite-time singularities of the gCLM model, and we refer the reader to [8, 14] for a more comprehensive survey on this subject.
For the original CLM model with $a=0$, finite-time singularity was established simultaneously with the proposal of the model in [10] by solving the equation explicitly with suitable initial data. It was only recently that Elgindi and Jeong [14] discovered an exact self-similar finite-time blowup for the CLM model. Based on this exact self-similar solution, they also proved in the same paper the existence of self-similar finite-time blowups from smooth initial data for $|a|$ small enough using a series expansion argument. Later, Elgindi, Ghoul, and Masmoudi [13] improved on this result by establishing the stability of those self-similar blowups for sufficiently small $|a|$. A similar result was also established independently in a work of Chen, Hou, and Huang [6]. In a similar spirit, Lushnikov, Silantyev, Siegel [21] and Chen [4] independently found an exact self-similar solution for $a=1/2$. Chen also proved stable self-similar finite-time singularities from smooth initial data for $a$ close to $1/2$ using the method developed in [6]. Finite-time singularity in the case $a=1$ was conjectured by De Gregorio [12] and was first rigorously established in a self-similar form by Chen, Hou, and Huang [6] using a computer-assisted proof (the same proof framework as in [7]). Later, Huang, Tong, and Wei [17] further showed that the De Gregorio model actually admits infinitely many self-similar finite-time blowup solutions of the same blowup scaling but with distinct profiles (under re-scaling) that all correspond to the eigen-functions of a self-adjoint, compact operator. All these self-similar finite-time blowup results above concern only the cases of $a=0,1/2,1$ and $a$ very close to these three values, where the (approximate) self-similar profiles were either found explicitly or computed numerically. Moreover, self-similar blowups with $C^{\alpha}$ profiles have also been constructed [14, 24] for a continuous range of $a$ under the constraint $|a|\alpha\ll 1$. Finally, Huang, Qin, Wang, and Wei [16] proved the existence of self-similar finite-time blowups with (interiorly) smooth profiles of the gCLM model for all $a\leq 1$ via a novel unified fixed-point approach.
The main difficulty of modifying our fixed-point method from the gCLM case to the HL case lies in that the gCLM model is an equation of one active scalar $\omega$, while the HL model is a coupled system of two scalars $\omega,\theta$. Critically and surprisingly, though the nonlinear map $\bm{R}$ for the HL system is formally much more complicated than the one for the gCLM equation, it still enjoys the crucial property that it preserves the previously mentioned monotonicity and convexity from $f$ to $\bm{R}(f)$. There are two major reasons why this could work. Firstly, from a more essential perspective, this nice property of $\bm{R}$ owes to the kernel structure of the Hilbert transform $\bm{H}$, which has already been exploited in [16]. Secondly, the coupling between $\omega$ and $\theta$ is not too complicated in the sense that the $\theta$ equation in (1.1) does not involve $\omega$ explicitly. Later, we will see that this fact enables us to construct the map $\bm{R}$ hierarchically with some intermediate maps (including the map $\bm{M}$ that is new in this work). It is worth mentioning that such coupling structure also appears in the CKY model [9, 8], which takes the same form as in (1.1) except that the relation $u_{x}=\bm{H}(\omega)$ is replaced with $(u/x)_{x}=\omega/x$. Therefore, it might also be possible to apply our fixed-point method to the CKY model and prove the existence of a self-similar finite-time blowup. Moreover, we believe that this fixed-point framework can be further generalized to prove the existence of an exact self-similar finite-time blowup of the 2D Boussinesq system (1.2) that has a similar coupling structure between $\omega$ and $\theta$. This shall be our next step in this line of research.
@@ 2.1. Self-similar profiles
$\begin{split}&-(T-t)^{c_{\omega}-1}c_{\omega}\Omega+(T-t)^{c_{\omega}-1}c_{l}X\Omega_{X}+(T-t)^{2c_{\omega}}U\Omega_{X}=(T-t)^{c_{\theta}-c_{l}}\Theta_{X},\\ &-(T-t)^{c_{\theta}-1}c_{\theta}\Theta+(T-t)^{c_{\theta}-1}c_{l}X\Theta_{X}+(T-t)^{c_{\theta}+c_{\omega}}U\Theta_{X}=0,\\ &U_{X}=\bm{H}(\Omega),\end{split}$
@@
It is worth mentioning that Chen, Hou, and Huang [7] studied the dynamic rescaling equations of the HL model (see [7, Equation (2.1)]),
@@
Finally, we remark that it remains open whether there exist nontrivial odd solutions of equations (2.1) that are more degenerate at $x=0$ in the sense that $\omega^{\prime}(0)=v^{\prime}(0)=0$ while $\omega^{(n)}(0)\neq 0,v^{(n)}(0)\neq 0$ for some odd integer $n>1$. Nevertheless, numerical evidence of more degenerate self-similar finite-time blowup solutions of the HL model has been reported in the thesis of Liu [20]. It is also worth noting that a novel multi-scale self-similar finite-time blowup phenomenon has recently been discovered and proved by Huang, Qin, and Wang [19] for the CLM model (i.e. (1.5) with $a=0$) for more degenerate initial data, which is qualitatively different from the classical one-scale self-similar blowup constructed in [14] with the non-degeneracy condition ($\omega^{\prime}(0)\neq 0$).
@@ 2.2. Reformulation of the problem
Note that the second equation of (2.8) does not involve $f$ explicitly. We then easily get
@@
meaning that $m$ can be explicitly determined by $g$. Next, we rearrange the first equation of (2.8) to get
@@
$f(x)=x^{-d}\exp\left((d-1)\int_{0}^{x}\frac{g(y)-1}{yg(y)}\,\mathrm{d}{y}\right)\cdot\int_{0}^{x}d\,y^{d-1}\frac{m(y)}{g(y)}\exp\left((d-1)\int_{0}^{y}\frac{1-g(z)}{zg(z)}\,\mathrm{d}{z}\right)\,\mathrm{d}{y}.$
@@
$\begin{split}f\quad\longrightarrow\quad&b=-\bm{H}(xf)(0),\quad c=c(f),\quad d=\frac{b+c}{2c},\\ \longrightarrow\quad&g=1-\frac{(-\Delta)^{-1/2}(xf)/x-b}{c},\\ \longrightarrow\quad&m=\frac{1}{g}\cdot\exp\left(2\int_{0}^{x}\frac{1-g}{yg}\,\mathrm{d}{y}\right),\\ \longrightarrow\quad&f=x^{-d}\exp\left((d-1)\int_{0}^{x}\frac{g-1}{yg}\,\mathrm{d}{y}\right)\cdot\int_{0}^{x}d\,y^{d-1}\frac{m}{g}\exp\left((d-1)\int_{0}^{y}\frac{1-g}{zg}\,\mathrm{d}{z}\right)\,\mathrm{d}{y}.\end{split}$ (2.9)
From the derivation above, it is apparent that a solution of the integral equations (2.9) corresponds to a solution of our main equations (2.1) via the change of variables (2.6), (2.7) (with $c_{\omega}$ given by the second equation in (2.5)).
@@ 3. Existence of solution by a fixed-point method
Our goal of this section is to show that the nonlinear integral equations (2.9) admit a non-trivial solution. As we can see, the system (2.9) naturally defines a fixed-point problem of a nonlinear nonlocal map. To prove the existence of a suitable fixed-point solution, we need to construct some appropriate function set in a Banach function space on which we can establish continuity and compactness of this nonlinear map, and then we use the Schauder fixed-point theorem (Fact B.1).
@@ 3.1. A fixed-point problem
The upper bound of $f$ in (3.2) follows from the assumptions that $f(\sqrt{s})$ is convex in $s$, $f^{\prime}_{-}(1)\leq-\eta$, and $f(x)$ is non-increasing on $0,+\infty)$.
We remark that, though a function $f\in\mathbb{D}$ is not required to be differentiable, the one-sided derivatives $f^{\prime}_{-}(x)$ and $f^{\prime}_{+}(x)$ are both well defined at every point $x$ by the convexity of $f(\sqrt{s})$ in $s$. In what follows, we will abuse notation and simply use $f^{\prime}(x)$ for $f^{\prime}_{-}(x)$ and $f^{\prime}_{+}(x)$ in both weak sense and strong sense. For example, when we write $f^{\prime}(x)\leq C$, we mean $f^{\prime}_{-}(x)\leq C$ and $f^{\prime}_{+}(x)\leq C$ at the same time. In this context, the non-increasing property of $f$ on $0,+\infty)$ can be represented as $f^{\prime}(x)\leq 0$ for $x\geq 0$.
@@
One can verify that $\bm{T}(f)(0)=0$ for $f\in\mathbb{D}\subset L^{1}(\mathbb{R})$. Comparing this definition with (2.2), we find
@@
Note that $\bm{M}(f)(0)=1/\bm{G}(f)(0)=1$ for $f\in\mathbb{D}$. Finally, letting
@@
We now explain the design of the set $\mathbb{D}$. The guiding idea is to make $\mathbb{D}$ as specific as possible while keeping it compact in the $L^{\infty}_{\rho}$-norm and closed under the map $\bm{R}$, so that we can apply the Schauder fixed-point theorem. The compactness of $\mathbb{D}$ is relatively easy, while its closedness under $\bm{R}$ takes most of the work. Firstly, the monotonicity of $f(x)$ in $x$ and the convexity of $f(\sqrt{s})$ in $s$ for $f\in\mathbb{D}$, which provide powerful controls on the functions in $\mathbb{D}$, are magically preserved by the map $\bm{R}$. This is essentially due to the nice properties of the kernel of the linear map $\bm{T}$. Secondly, the monotonicity, the convexity, and some calculations of $\bm{T}$ will lead to the control $m_{1}(x)\leq\bm{R}(f)\leq 1$. Hence, we can impose this condition on $\mathbb{D}$ as well. Thirdly, we need $\mathbb{D}\subset L^{1}(\mathbb{R})$ in order for $b(f)$ to be well defined on $\mathbb{D}$, which explains the condition $f(x)\lesssim|x|^{-1-\delta_{1}}$. However, to make sure this tail bound is not compromised by $\bm{R}$, we need the weaker estimate $f(x)\lesssim|x|^{-\delta_{0}}$, and to further pass this weaker estimate to $\bm{R}(f)$ we need to use the condition $f^{\prime}_{-}(1)\leq-\eta$. Also, $f^{\prime}_{-}(1)\leq-\eta$ leads to the upper bound in (3.2), which further implies $c(f)>0$, so that $\bm{G}(f)$ and $d(f)$ are both well-defined. Finally, the estimate $\bm{R}(f)^{\prime}(1)\leq-\eta$ can be derived from $f^{\prime}_{-}(1)\leq-\eta$ and the convexity of $f(x)$ in $x^{2}$, therefore closing the loop.